OpenStudy (anonymous):

A circle is shown on the coordinate plane below. What is the equation of this circle?

OpenStudy (anonymous):

x2 + (y + 1)2 = 4 x2 + (y – 1)2 = 2 x2 + (y + 1)2 = 2 x2 + (y – 1)2 = 4

OpenStudy (anonymous):

x2 + (y – 1)2 = 4

OpenStudy (anonymous):

Thnx Bro

OpenStudy (anonymous):

welcome just simple use\[ (x-h)^2+(y-k)^2=r^2\]